# From Reliability to Security: How RIS-Assisted Adaptive SM and SSK Enhances Wireless Systems

### v1

@@ III Generalized Pre-Defined ACs Selection
Since the number of available ACs must be a power of 2 and is more than what is required, e.g., 3 receive antennas lead to 7 ACs but 4 are available to represent the transmitted bit sequences, we need the generalized pre-defined ACs selection to eliminate the unnecessary ACs. Before the transmission, the transmitter and receiver have an agreement that only $D$ ACs can be the indices to convey the bits sequence, where $D=\lfloor\log_{2}\left(N_{r}-1\right)\rfloor$ and $D$ ACs are selected by the designed generalized pre-defined ACs selection. These $D$ ACs are selected through the specific order. [^3] Our aim in this paper is primarily establishing and showing a fundamental system designs and comprehensive performance analysis. The discussion for this extension are out of scope and space limitations, as the potential investigation in future works. In our proposed scheme, we eliminate the AC with the maximum number of antennas. For instance, when $N_{t}=4$, we will remove the ACs with $N_{a}=4$. Then, we sort the remaining ACs in ascending order based on $N_{a}$. The first $D$ ACs are selected for bit mapping. By doing so, we can not only remove the ACs with the maximum number of $N_{a}$ but also minimize the involvement of ACs with higher $N_{a}$ in subsequent signal processing stages, which can further reduce the effects from the non-constructive part. For example, we have $N_{t}=4$ and select 8 ACs based on this order from the 15 available ACs through the aforementioned steps. In this way, the AC with $N_{a}=4$ and most part of ACs with $N_{a}=3$ are eliminated. After this, the AC set can be given as $\varLambda(r)$ with $r=1,2,\cdots,D$. The designed selection can significantly reduce the detection complexity, and a detailed analysis of its detection complexity is provided in sub-section B of Sec. IV. The generalized pre-defined ACs selection works at the intelligent unit of the RIS after the information of selected ACs inputting. Then, the intelligent unit controls various parts of REs to amplify the transmitted signal for the specific receive antenna/antennas.
@@ IV-B1 ABER for the RASM Scheme
The analytical ABER performance of the RASM scheme is given. According to Eq. (8) of ML detector, the instantaneous pairwise error probability (PEP) can be expressed by giving channel $\mathrm{h}_{1}^{i}$ and $\mathrm{h}_{2}^{i,l}$ as
@@
which applies the moment generating function (MGF) of $Z$, given as $M_{Z}\left(t\right)=\int_{0}^{\infty}{e^{tZ}dZ}$, with $t=\frac{-1}{4\sin^{2}\!\>\tau N_{0}}$. Actually, $Z$ is the squared Euclidean distance (SED) between the symbols $G_{r,k}$ and $G_{\hat{r},\hat{k}}$. Here, the MGF of Eq. (32) can be derived by considering the general quadratic form of correlated Gaussian random variables and counts on erroneous or correct detection of the $r$-th and $\hat{r}-$th AC indices. Note that the error in the detection of AC index is able to affect the accuracy of constellation symbol detection. Thus, we separate $Z$ into two categories, which are given as: $Z_{1}=\left\{Z\mid r\neq\hat{r}\right\}$ and $Z_{2}=\left\{Z\mid r=\hat{r}\right\}$.
@@
where $q_{1}$ and $q_{2}$ follow complex Gaussian distribution by applying central limit theorem (CLT).  [^4] Note that $\alpha$ and $\beta$ are the magnitudes of standard complex Gaussian random variables, they follow the Rayleigh distribution with mean value $\frac{\sqrt{\pi}}{2}$ variance $\frac{4-\pi}{4}$, respectively. Moreover, since $\omega_{i}$, $\varphi_{i}$, and $\theta_{i}$ are the phase of standard complex Gaussian random variable and follow the uniform distribution at the range in $(-\pi,\pi)$. The mean vector $\boldsymbol{\mu}_{1}$ and covariance matrix $\boldsymbol{\sigma}_{1}^{2}$ of $\boldsymbol{z}_{1}=\left[\left(q_{1}\right)_{\mathfrak{R}},\left(q_{1}\right)_{\mathfrak{T}},\left(q_{2}\right)_{\mathfrak{R}},\left(q_{2}\right)_{\mathfrak{T}}\mathrm{}\right]$, which are respectively given as follows:
@@ V-A System Model with Eavesdropper
Since the adaptive antenna selection works at a specific receiver in both proposed schemes, the eavesdropper cannot completely detect and decode the received signals from the RIS, which further indicates that the proposed schemes offer better secrecy performance in terms of PLS during transmission. A new system model in Fig. 2 includes an additional eavesdropper, referred to as Eve equipped with one antenna,  [^6] We primarily emphasize the superiority of PLS performance in our proposed schemes through comparisons with related benchmarks. Accordingly, Eve is assumed to possess equivalent radio hardware to Bob, e.g., single RF chain, identical noise variance, but lacks Bob’s private seed for RIS reconfiguration. Eve can access statistical channel parameters for transmitter-to-RIS and RIS-to-receiver links, and measure her own instantaneous realizations via public pilots. However, scenarios involving Eve’s physical access to the RIS controller or active jamming, e.g., active eavesdropping or jamming attacks, represent different threat models beyond our scope and are not considered. which aims at eavesdropping the transmitted signals from the RIS to the receiver as Bob. It is worth noting that this system model in Fig. 2 is suitable for both the RASM and RASSK schemes, with the only difference that the constellation symbols are transmitted in the RASM scheme instead of the RASSK one. Besides, Eve is assumed to be passive and placed as close to Bob as possible, which means Eve cannot actively attack or interrupt. Here, we also assume that Eve has perfect knowledge of the CSI from the RIS to the receive antenna and the transmitter to the RIS which is given as $g_{2}\sim\mathcal{C}\mathcal{N}\left(0,1\right)$, following Rayleigh distribution [26]-[28]. Thus, the received signals at Eve can be given as
@@
where $g_{2}^{i}$ and $h_{1}^{i}$ bit sequence $i$-th RE respectively, $\Phi_{i}$ stands for the adjusted phase from the $i$-th RE, and $n_{e}$ is the white Gaussian noise with variance $N_{0}$. The ML detector at Eve is
@@ V-C Secrecy Rate Analysis for the RASSK Scheme
VI Simulation Results
In this section, the analytical and simulation results are presented. To provide a more comprehensive illustration of the performance in the proposed scheme, a uniform assumption of Rayleigh channel fading is adopted, excepted for Fig. 6, with the transmission antenna $N_{t}=1$ to be considered in all cases of the RASM and RASSK schemes. All of these results are conducted by employing the Monte Carlo simulations, encompassing a substantial number of channel realizations, precisely $3\times 10^{6}$. Similar to [10]-[14], $\bar{\gamma}=\frac{E_{s}}{N_{0}}$ is considered as the average SNR at receiver.
The analytical and simulation results of BER in the RASM scheme are presented in Fig. 3, where the results of the cases with $N_{r}=$ 4 and 5, N=N=8 and 16, M=M=2 and 4 are given, respectively. As depicted in Fig. 3, BER decreases monotonically with an increase in $N$. This trend underscores the significant impact of the RIS in enhancing BER performance. For the case with $N_{r}=4$, $N=16$, and $M=2$, we can find that the BER of the RASM scheme is lower than $10^{-5}$ when SNR is equal to 10 dB and bpcu is equal to 4, which indicates the proposed scheme can achieve the desired BER performance. Particularly as the SNR increases. the simulation results closely approach to the theoretical ones, which is mainly because of the applications of CLT in theoretical analysis. Additionally, the gap between the theoretical and simulation results decreases as $N$ increases. This is also because the CLT utilization, causing the theoretical BER to converge with the simulations as $N$ grows. Moreover, we can observe that the BER performance worsens as $N_{r}$ increases, which is because the increase in $N_{r}$ leads to the increase in number of ACs, thus leading to increase error possibility in detection. For the same reason, by comparing the results of $M=2$ and $M=4$ between (a) and (b) in Fig. 3, the slight discrepancy and learn that the BER performance deteriorates with the increase of $M$. Fig. 3 (c) illustrates the analytical and simulation BER performance of the RASSK scheme, where $N_{r}=$3, 5, 6 and N=N=8, 16, respectively. Similar to (a) and (b), we can learn from (c) that BER performance of the RASSK scheme gets better with the increase of SNR and the analytical BER is approached to the simulation one especially in the higher SNR range. Additionally, the results in (c) also point out that the RASSK has satisfactory BER performance even under the low SNR situation.
@@ VI Simulation Results
Giving the same condition on SE, the comparison of the various received RIS-assisted IM schemes with the imperfect CSI estimation is shown in Fig. 5, where the schemes include the RASM, RASSK, and RSM schemes. Here, the imperfect CSI is assumed in $h_{2}^{i,m}$ with $\delta^{2}$ in the RTSM and RTSSK schemes representing the variance of estimation error [31]-[32] and the spectral efficiency is R=R=4 bpcu. Detail explanations of the imperfect CSI estimation can be found in Appendix B. Results in Fig. 5 show that the robustness of the RASSK scheme is better than that of the RASM and RSM schemes, despite the fact that the $\delta^{2}$ in the RASSK results is larger than others. Also, the BER performance of RASM scheme outperforms the one of the RSM scheme. According to the comparisons in Fig. 5, we can find that the proposed RASSK scheme has more potential to apply in complex wireless environment, due to its better robustness.
Fig. 6 shows the BER performance of RASM scheme with $N_{r}=8$$N_{d}=4$,$N=32$, and $M=2$, RASSK scheme with $N_{r}=8$ and $N_{d}=5$, RSM scheme with $N_{r}=8$, $N=32$, and $M=2$, of which the K=K=5, 10, and 15, respectively. In Fig. 6, we can find that the BER performance of all schemes gets worse with the increase of $K$. This is because stronger LoS component makes the channel responses at different antennas more similar. When receive antennas are closely spaced, their channel characteristics become very similar, thus reducing spatial diversity. Since IM schemes, especially the SM and SSK related schemes, rely on distinct channel characteristics for correct detection of antenna indices, this reduced diversity can lead to degrade the BER performance. This phenomenon is particularly evident in the results of RASSK scheme shown in Fig.6, which further corroborates the aforementioned argument. However, the BER performance of RASM scheme still outperforms the one of RSM scheme even with different $K$.
Figure 5: BER performance of the RASM, RASSK, and RASM schemes with imperfect CSI estimation.
@@
Following model of MRC in [33], the SR of the RIS-assisted SIMO scheme with MRC (SIMO-MRC), the TSM scheme with MRC (TSM-MRC), the TSM-MRC with AN  [^9] This AN is designed with the noise variance $N_{0}$ as the same to the Gaussian white noise. Detail AN model can be found in [34], which is not presented in this paper., and the TSSK scheme with MRC (TSSK-MRC) are given to compare with the one in both proposed schemes as shown in Fig. 7.101010Since there is currently no existing literature discussing and investigating the PLS performance of the TASM scheme, where the adaptive antenna selection works at transmitter and the RIS is employed as a passive relay, we thus select the TSM and TSSK schemes as benchmarks to compare the SR with both proposed schemes in this paper. To make a fair comparison, as the same SE in 3 bits/s/Hz, the RASM has $N_{r}=$4, N=N=8, M=M=2, and the RASSK has $N_{r}=$4, N=N=8, and the SIMO-MRC has $N_{t}=$1 $N_{r}=$3, N=N=8, M=M=32, and the TSM-MRC scheme has $N_{t}=$3 $N_{r}=$3, N=N=8, M=M=2, and the TSSK-MRC scheme has $N_{t}=$8 $N_{r}=$3, N=N=8. In Fig. 7, the SR of both proposed schemes is higher than that of the others from -30 dB to 30 dB, demonstrating that the received IM schemes achieve better PLS performance compared to the traditional SIMO-MRC, TSM-MRC, and TSSK-MRC schemes, even the PLS wireless technique, e.g., AN in the TSM-MRC. Below around -5 dB, the SR of the SIMO-MRC scheme is slightly higher than that of TSM-MRC and TSSK-MRC, indicating that the PLS performance based on transmit antenna indices is weaker than that based on constellation symbols. The SR results of TSM-MRC and TSSK-MRC are closed because Bob and Eve share the same transmitter-to-RIS channel, and the CSI is assumed to be known to Eve, allowing her to fully detect and decode the transmitted signals, which is a limitation of transmitted IM schemes. Notably, the SR of the SIMO-MRC, TSM-MRC, and TSSK-MRC schemes approaches but does not reach zero, as Bob’s channel capacity remains higher than Eve’s due to having multiple antennas, while Eve only has one.

### v2

@@ II-C RASSK Transmission
II-D Impact of Quantized RIS Phase Errors
In practical deployments, achieving perfect continuous phase shifts is highly challenging due to hardware impairments and the limited resolution of phase shifters, i.e., finite discrete phase quantization. To account for these practical hardware limitations, a phase error term $\Delta\varphi_{i,m}$ can be introduced. The ideal adjusted phase matrix $\mathbf{\Phi}_{r}$ is replaced by the practical phase matrix with errors, denoted as $\tilde{\mathbf{\Phi}}_{r}$, which is given as
$\tilde{\mathbf{\Phi}}_{r}=\text{diag}\left(e^{j(\varphi_{1,m}+\Delta\varphi_{1,m})},\dots,e^{j(\varphi_{N,m}+\Delta\varphi_{N,m})}\right),$ (17)
with $\Delta\varphi_{i,m}$ representing the phase error of the $i$-th reflecting element intended for the $m$-th receive antenna. For a $b$-bit quantized RIS, this phase error is typically modeled as an independent uniform random variable distributed over $\frac{\pi}{2^{b}},\frac{\pi}{2^{b}}]$. Consequently, the received signal expressions for the RASM and RASSK schemes, originally given in Eq. (2) and Eq. (10), respectively, must be modified to incorporate these hardware impairments:
$\tilde{\boldsymbol{Y}}_{\text{RASM}}=\mathbf{H}_{2}\tilde{\mathbf{\Phi}}_{r}\mathbf{H}_{1}x_{k}+\mathbf{N},$ (18)
$\tilde{\boldsymbol{Y}}_{\text{RASSK}}=E_{s}\mathbf{H}_{2}\tilde{\mathbf{\Phi}}_{r}\mathbf{H}_{1}+\mathbf{N}.$ (19)
Due to the presence of the phase error $\Delta\varphi_{i,m}$, the perfect passive beamforming condition ,e.g., $\varphi_{i}=\omega_{i,l}+\theta_{i}$, required to maximize the SNR cannot be strictly satisfied.  [^4] While this foundational work evaluates the proposed schemes under imperfect CSI and Rician fading, practical deployments face additional impairments like CSI feedback delays, hardware non-idealities, and frequency-selective fading. Incorporating these dynamic factors would render the closed-form analysis mathematically intractable. Therefore, a comprehensive investigation of these practical impairments is left for future research.
@@ III Generalized Pre-Defined ACs Selection
Since the total number of mathematically available ACs is $J=2^{N_{r}}-1$, but the number of required ACs for bit mapping must be a power of 2, the available ACs are typically more than required. For instance, 3 receive antennas yield 7 non-empty ACs, but only 4 are required to represent 2 bits. Therefore, a generalized pre-defined ACs selection algorithm is essential to eliminate the unnecessary ACs. Before transmission, the transmitter and the receiver pre-agree that exactly $D$ ACs will be utilized to convey the $b_{1}=\lfloor\log_{2}(2^{N_{r}}-1)\rfloor$ bits, where $D=2^{b_{1}}$. These $D$ ACs are chosen from the set of non-empty receive subsets through a specific sequence. In our proposed scheme, we first systematically eliminate the ACs containing the maximum number of antennas, i.e., $N_{a}=N_{r}$. Subsequently, we sort the remaining ACs in ascending order based on their $N_{a}$. Finally, the first $D$ ACs in this sorted list are selected for bit mapping. By employing this method, we not only exclude the ACs with the maximum $N_{a}$ but also minimize the involvement of ACs with higher $N_{a}$ in subsequent signal processing stages, which effectively mitigates the impacts of the non-constructive interference. As a concrete example, assuming $N_{r}=4$, we need to select $D=2^{3}=8$ ACs from the 15 available ones. By following the aforementioned steps, the single AC with $N_{a}=4$ and a majority of the ACs with $N_{a}=3$ are systematically eliminated. The final selected AC set is denoted as $\Lambda(r)$ with $r=1,2,\dots,D$. This designed selection significantly reduces the detection complexity, as detailed in Sec. IV.C. The designed selection can significantly reduce the detection complexity, and a detailed analysis of its detection complexity is provided in sub-section B of Sec. IV. The generalized pre-defined ACs selection works at the intelligent unit of the RIS after the information of selected ACs inputting. Then, the intelligent unit controls various parts of REs to amplify the transmitted signal for the specific receive antenna/antennas. By intelligently pruning the transmission states rather than relying solely on post-processing at the receiver, this mechanism intrinsically resolves the complexity bottleneck of dynamic RIS beamforming.
@@ IV-B1 ABER for the RASM Scheme
The analytical ABER of the RASM scheme is given. According to Eq. (8) of ML detector, the instantaneous pairwise error probability (PEP) can be expressed by giving channel $\mathrm{h}_{1}^{i}$ and $\mathrm{h}_{2}^{i,l}$ as
@@
which applies the moment generating function (MGF) of $Z$, given as $M_{Z}\left(t\right)=\int_{0}^{\infty}{e^{tZ}dZ}$, with $t=\frac{-1}{4\sin^{2}\!\>\tau N_{0}}$. Actually, $Z$ is the squared Euclidean distance (SED) between the symbols $G_{r,k}$ and $G_{\hat{r},\hat{k}}$. Here, the MGF of Eq. (35) is derived by the general quadratic form of correlated Gaussian random variables and counts on erroneous or correct detection of the $r$-th and $\hat{r}-$th AC indices. Note that the error in the detection of AC index is able to affect the accuracy of constellation symbol detection. Thus, we separate $Z$ into two categories, which are given as: $Z_{1}=\left\{Z\mid r\neq\hat{r}\right\}$ and $Z_{2}=\left\{Z\mid r=\hat{r}\right\}$.
@@
where $q_{1}$ and $q_{2}$ follow complex Gaussian distribution by applying central limit theorem (CLT).  [^5] Note that $\alpha$ and $\beta$ are the magnitudes of standard complex Gaussian random variables, they follow the Rayleigh distribution with mean value $\frac{\sqrt{\pi}}{2}$ variance $\frac{4-\pi}{4}$, respectively. Moreover, since $\omega_{i}$, $\varphi_{i}$, and $\theta_{i}$ are the phase of standard complex Gaussian random variable and follow the uniform distribution at the range in $(-\pi,\pi)$.  [^6] It should be noted that the derived ABER expressions rely on the CLT approximation, which is asymptotically accurate. In practical RIS-assisted systems, the reliability of this approximation depends on the number of reflecting elements $L_{i}$ assigned to each receive antenna. Through numerical validation, we observe that the approximation remains highly precise when $L_{i}\geq 8$. As $L_{i}$ or the total RIS size $N$ increases, the aggregate channel distribution further converges to the complex Gaussian model, making the analytical results more precise. The mean vector $\boldsymbol{\mu}_{1}$ and covariance matrix $\boldsymbol{\sigma}_{1}^{2}$ of $\boldsymbol{z}_{1}=\left[\left(q_{1}\right)_{\mathfrak{R}},\left(q_{1}\right)_{\mathfrak{T}},\left(q_{2}\right)_{\mathfrak{R}},\left(q_{2}\right)_{\mathfrak{T}}\mathrm{}\right]$, which are respectively given:
@@ IV-B2 ABER for the RASSK Scheme
Remark 3
The theoretical rigor of this framework is established as follows: First, the ABER expressions in Eq. (56) and Eq. (59) provide strict theoretical upper bounds derived via the MGF approach. Second, the deterministic even division strategy for RIS phase adjustment circumvents the need for algorithmic convergence proofs, while the ML detector ensures detection optimality. Although employing the CLT for mathematical tractability, the tightness of this asymptotic analysis is robustly validated by Monte Carlo simulations in Sec. VI.
@@ IV-C Detection Complexity
IV-D Energy Efficiency Analysis
In addition to SE and BER, energy efficiency (EE) is a paramount metric for evaluating the sustainability of modern wireless communication systems, particularly in IoT scenarios. To mathematically formulate the EE for the proposed schemes, we first define the total power consumption model, denoted as $P_{total}$. Unlike traditional active MIMO systems that require multiple power-hungry active RF chains at the transmitter, the proposed schemes rely on a single transmit RF chain and a passive RIS array. Therefore, $P_{total}$ can be modeled as:
$P_{total}=\frac{P_{t}}{\eta}+P_{TX}+N_{r}P_{RX}+NP_{RE}+P_{RIS\_ctrl},$ (62)
where $P_{t}$ is the actual transmit power, and $\eta$ represents the efficiency of the power amplifier and $P_{TX}$ is the hardware static power consumed by the single RF chain at the transmitter. As all $N_{r}$ receive antennas remain active to harvest both constructive and non-constructive signals in our schemes, $P_{RX}$ denotes the power consumption per active receive RF chain. For the RIS, $P_{RE}$ represents the near-zero power consumed by each of the $N$ passive reflecting elements for phase shifting, and $P_{RIS\_ctrl}$ is the power consumption of the intelligent controller unit. Thus, the EE for the RASM and RASSK schemes can be directly derived by substituting their respective SE from Eq. (21) and Eq. (22) into the EE definition:
$EE_{RASM}=\frac{\lfloor\log_{2}(2^{N_{r}}-1)\rfloor+\log_{2}M}{\frac{P_{t}}{\eta}+P_{TX}+N_{r}P_{RX}+NP_{RE}+P_{RIS\_ctrl}},$ (63)
$EE_{RASSK}=\frac{\lfloor\log_{2}(2^{N_{r}}-1)\rfloor}{\frac{P_{t}}{\eta}+P_{TX}+N_{r}P_{RX}+NP_{RE}+P_{RIS\_ctrl}}.$ (64)
Remark 7
It is critical to highlight that the power consumption of a passive RIS element is significantly lower than that of an active RF chain ($P_{RE}\ll P_{TX},P_{RX}$). By leveraging the RIS to execute adaptive spatial modulation at the receiver side, the RASM and RASSK schemes effectively map extra information bits without activating additional transmit RF chains. Therefore, the numerator (SE) is substantially enhanced while the denominator ($P_{total}$) only experiences a marginal increase from the low-power RIS elements. This architectural advantage guarantees that the proposed schemes can achieve superior EE compared to conventional multiple-active-antenna transmission architectures.
@@ V-A System Model with Eavesdropper
Since the adaptive antenna selection works at a specific receiver in both proposed schemes, the eavesdropper cannot completely detect and decode the received signals from the RIS, which further indicates that the proposed schemes offer better secrecy performance in terms of PLS during transmission. A new system model in Fig. 2 includes an additional eavesdropper, referred to as Eve equipped with one antenna,  [^8] Scenarios involving Eve’s physical access to the RIS controller or active jamming, e.g., active eavesdropping or jamming attacks, represent different threat models beyond our scope and are not considered. which aims at eavesdropping the transmitted signals from the RIS to the receiver as Bob. It is worth noting that this system model in Fig. 2 is suitable for both the RASM and RASSK schemes, with the only difference that the constellation symbols are transmitted in the RASM scheme instead of the RASSK one. Besides, Eve is assumed to be passive and placed as close to Bob as possible, which means Eve cannot actively attack or interrupt. Here, we also assume that Eve has perfect knowledge of the CSI from the RIS to the receive antenna and the transmitter to the RIS which is given as $g_{2}\sim\mathcal{C}\mathcal{N}\left(0,1\right)$, following Rayleigh distribution [33]-[35]. Thus, the received signals at Eve can be given as
@@
where $g_{2}^{i}$ and $h_{1}^{i}$ bit sequence $i$-th RE respectively, $\Phi_{i}$ stands for the adjusted phase from the $i$-th RE, and $n_{e}$ is the white Gaussian noise with variance $N_{0}$.  [^9] Since $\Phi_{i}$ is optimized exclusively for Bob’s instantaneous channel, it is statistically independent of Eve’s channel $g_{2}^{i}$ assuming spatial separation. Consequently, $\Phi_{i}$ acts as a random phase value for Eve, precluding any coherent beamforming gain in subsequent derivations. The ML detector at Eve is
@@ V-C Secrecy Rate Analysis for the RASSK Scheme
Remark 8
Our PLS analysis is rooted in strict information-theoretic secrecy, not obfuscation. Following Kerckhoffs’s principle [38, 39], we assume Eve possesses full system knowledge, including the alphabet $M$, mapping table $\Lambda$, and RIS strategy. Security fundamentally stems from spatial channel decorrelation: the RIS passive beamforming is exclusively aligned to the legitimate cascaded channel. Consequently, an eavesdropper at a different location experiences an independent, unaligned channel. This physical disparity inherently degrades Eve’s mutual information and guarantees a positive secrecy rate, completely independent of her knowledge of the mapping rules.
Remark 9
The RIS phase shifts are strictly optimized for the legitimate cascaded channel. Consequently, even if an eavesdropper intercepts high signal energy, their independent cascaded channel scrambles the delicate energy differences and phase alignments required to decode the spatial activation cues, rendering statistical blind estimation ineffective. This evaluation assumes the ubiquitous condition of independent fading, e.g., $d\gg\lambda/2$. However, if the eavesdropper is located in extreme proximity to the legitimate receiver ($d<\lambda/2$), their channels become highly spatially correlated. Under such specific conditions, significant information leakage is inevitablea fundamental physical limitation governed by electromagnetic wave propagation that is intrinsic to all spatial-domain physical layer security architectures.
@@
VI Simulation Results
In this section, the analytical and simulation results are presented. To provide a more comprehensive illustration of the performance in the proposed scheme, a uniform assumption of Rayleigh channel fading is adopted, except for Fig. 6, with the transmission antenna $N_{t}=1$ to be considered in all cases of the RASM and RASSK schemes. All of these results are conducted by employing the Monte Carlo simulations, encompassing a substantial number of channel realizations, precisely $3\times 10^{6}$. Similar to [14]-[18], $\bar{\gamma}=\frac{E_{s}}{N_{0}}$ is considered as the average SNR at receiver.
The analytical and simulation results of BER in the RASM scheme are presented in Fig. 3, where the results of the cases with $N_{r}=$ 4 and 5, N=N=8 and 16, M=M=2 and 4 are given, respectively. As depicted in Fig. 3, BER decreases monotonically with an increase in $N$. This trend underscores the significant impact of the RIS in enhancing BER performance. For the case with $N_{r}=4$, $N=16$, and $M=2$, we can find that the BER of the RASM scheme is lower than $10^{-5}$ when SNR is equal to 10 dB and bpcu is equal to 4, which indicates the proposed scheme can achieve the desired BER performance.  [^11] Since the analytical curves represent theoretical upper bounds derived via the Union Bound technique, these bounds naturally become loose and may mathematically exceed 0.5 in the low-SNR regime due to the summation of overlapping pairwise error probabilities. Particularly as the SNR increases. the simulation results closely approach to the theoretical ones, which is mainly because of the applications of CLT in theoretical analysis. Additionally, the gap between the theoretical and simulation results decreases as $N$ increases. This is also because the CLT utilization, causing the theoretical BER to converge with the simulations as $N$ grows. We can observe that the BER performance worsens as $N_{r}$ increases, which is because the increase in $N_{r}$ leads to the increase in number of ACs, thus leading to increase error possibility in detection. For the same reason, by comparing the results of $M=2$ and $M=4$ between (a) and (b) in Fig. 3, the slight discrepancy and learn that the BER performance deteriorates with the increase of $M$. Fig. 3 (c) illustrates the analytical and simulation BER performance of the RASSK scheme, where $N_{r}=$3, 5, 6 and N=N=8, 16, respectively. Similar to (a) and (b), we learn from (c) that BER performance of the RASSK scheme gets better with the increase of SNR and the analytical BER is approached to the simulation one especially in the higher SNR range. Additionally, the results in (c) also point out that the RASSK has satisfactory BER performance even under the low SNR situation.
@@ VI Simulation Results
Giving the same condition on SE, the comparison of the various received RIS-assisted IM schemes with the imperfect CSI estimation is shown in Fig. 5, where the schemes include the RASM, RASSK, and RSM schemes. Here, the imperfect CSI is assumed in $h_{2}^{i,m}$ with $\delta^{2}$ in the RTSM and RTSSK schemes representing the variance of estimation error [40]-[41] and the spectral efficiency is R=R=4 bpcu. Detail explanations of the imperfect CSI estimation can be found in Appendix B. Results in Fig. 5 show that the robustness of the RASSK scheme is better than that of the RASM and RSM schemes, despite the fact that the $\delta^{2}$ in the RASSK results is larger than others. Also, the BER performance of RASM scheme outperforms the one of the RSM scheme. Furthermore, when facing hardware-induced phase quantization errors ($b=3$), the RASSK scheme exhibits exceptional resilience, with its BER continuously dropping below $10^{-3}$ at low SNR regions. In contrast, the RASM scheme experiences a noticeable performance floor. This performance gap physically reveals that conveying information purely through received power at selected antennas (as in RASSK) is inherently more robust to phase mismatches than detecting conventional $M$-ary constellation symbols (as in RASM). According to these comprehensive comparisons, it can be firmly concluded that the proposed RASSK scheme holds immense potential for applications in complex and imperfect real-world wireless environments due to its superior robustness against both CSI estimation inaccuracies and hardware phase quantization errors.
Fig. 6 shows BER of RASM scheme with $N_{r}=8$$N_{d}=4$,$N=32$, and $M=2$, RASSK scheme with $N_{r}=8$ and $N_{d}=5$, RSM scheme with $N_{r}=8$, $N=32$, and $M=2$, of which the K=K=5, 10, and 15, respectively. In Fig. 6, we can find that the BER performance of all schemes gets worse with the increase of $K$. This is because stronger LoS component makes the channel responses at different antennas more similar. When receive antennas are closely spaced, their channel characteristics become very similar, thus reducing spatial diversity. Since IM schemes, especially the SM and SSK related schemes, rely on distinct channel characteristics for correct detection of antenna indices, this reduced diversity leads to degrade BER performance. This phenomenon is particularly evident in the results of RASSK scheme shown in Fig. 6, which further corroborates the aforementioned argument. However, the BER performance of RASM scheme still outperforms the one of RSM scheme even with different $K$. As shown in Fig. 6, RASM consistently outperforms RSM across varying $K$. This stems from RASM’s adaptive antenna activation, which generates richer spatial signatures and allows a lower constellation order, thereby increasing the minimum Euclidean distance. Also, the RIS phases are optimized based on the instantaneous cascaded CSI, which includes both LoS and Rayleigh components, constructively harvesting Rayleigh energy. Optimizing solely for the LoS path would leave the unaligned Rayleigh component as severe interference.
Figure 5: BER performance of the RASM, RASSK, and RASM schemes with imperfect CSI estimation and quantized RIS phase errors.
@@
Following model of MRC in [42], the SR of the RIS-assisted SIMO scheme with MRC (SIMO-MRC), the TSM scheme with MRC (TSM-MRC), the TSM-MRC with AN  [^13] This AN is designed with the noise variance $N_{0}$ as the same to the Gaussian white noise. Detail AN model can be found in [43]., and the TSSK scheme with MRC (TSSK-MRC) are given to compare with the one in both proposed schemes as shown in Fig. 7.141414Since there is currently no existing literature discussing and investigating the PLS performance of the TASM scheme, where the adaptive antenna selection works at transmitter and the RIS is employed as a passive relay, we thus select the TSM and TSSK schemes as benchmarks to compare the SR with both proposed schemes in this paper. To make a fair comparison, as the same SE in 3 bits/s/Hz, the RASM has $N_{r}=$4, N=N=8, M=M=2, and the RASSK has $N_{r}=$4, N=N=8, and the SIMO-MRC has $N_{t}=$1 $N_{r}=$3, N=N=8, M=M=32, and the TSM-MRC scheme has $N_{t}=$3 $N_{r}=$3, N=N=8, M=M=2, and the TSSK-MRC scheme has $N_{t}=$8 $N_{r}=$3, N=N=8. In Fig. 7, the SR of both proposed schemes is higher than that of the others from -30 dB to 30 dB, demonstrating that the received IM schemes achieve better PLS performance compared to the traditional SIMO-MRC, TSM-MRC, and TSSK-MRC schemes, even the PLS wireless technique, e.g., AN in the TSM-MRC. Below around -5 dB, the SR of the SIMO-MRC scheme is slightly higher than that of TSM-MRC and TSSK-MRC, indicating that the PLS performance based on transmit antenna indices is weaker than that based on constellation symbols. The SR results of TSM-MRC and TSSK-MRC are closed because Bob and Eve share the same transmitter-to-RIS channel, and the CSI is assumed to be known to Eve, allowing her to fully detect and decode the transmitted signals, which is a limitation of transmitted IM schemes. Notably, the SR of the SIMO-MRC, TSM-MRC, and TSSK-MRC schemes approaches but does not reach zero, as Bob’s channel capacity remains higher than Eve’s due to having multiple antennas, while Eve only has one.
Remark 10
Achieving symbol-level RIS reconfiguration demands strict synchronization, as timing failures can cause severe phase misalignment and substantial BER degradation, which is a non-trivial issue and out of our scope, thus leaving for future works. Regarding hardware complexity, both the proposed RASM and conventional fixed-antenna schemes, e.g., RSM, RGSM, and RGSSK, fundamentally require all $N_{r}$ receive antennas to be connected to active RF chains for spatial detection. Therefore, our adaptive mechanism introduces no additional RF hardware cost at the receiver. This receiver-centric complexity is a deliberate design trade-off, enabling a minimalist, single-RF-chain architecture for power-constrained IoT transmitters.